You are choosing between 2 plans at a discount warehouse. Plan A offers an annual membership fee of $300 and you pay 70% of the manufacturer's recommended list price. Plan B offers an anual membership fee of $40 and you pay 90% of the manufacturer's recommended list price. How many dollars of merchandise would have to purchase in a year to pay the same amount under both plans? What will be the cost for each plan?
Appreciate any help you can give me. Thanks

Answer by stanbon(75887)

You are choosing between 2 plans at a discount warehouse.
Plan A offers an annual membership fee of $300 and you pay 70% of the manufacturer's recommended list price.
A(p) = 300 + 0.70p
--------------------------
Plan B offers an annual membership fee of $40 and you pay 90% of the manufacturer's recommended list price.
B(p) = 40 + 0.90p
----------------------------
How many dollars of merchandise would have to purchase in a year to pay the same amount under both plans? What will be the cost for each plan?
-----
Equation:
300 + 0.7p = 40 + 0.9p
260 = 0.2p
p = $1300
====================================
Cheers,
Stan H.